peanuts sell for $3 per pound. cashews sell for $6 per pound. how many pounds of cashews should be mixed with 12 pounds of peanuts to obtain a mixture that sells for $4.20 per pound?
Answer by checkley77(12844)

12*3+6x=4.20(12+x)
36+6x=50.4+4.2x
6x-4.2x=50.4-36
1.8x=14.4
x=14.4/1.8
x=8 pounds of cashews is used.
Proof:
12*3+6*8=4.2(12+8)
36+48=4.2*20
84=84